The citizens of Exland, whose currency is the bun, conduct the transactions
outlined in the table.
Item Billions of buns
Imports of goods and services 600
Exports of goods and services 700
Foreign investment in Exland 350
Exland's investment abroad 430
Increase in the official holdings of foreign currency 20
(a) What is Exland's current account balance?
(b) What is Exland's capital account balance?
(c) What is Exland's official settlement balance?
(d) What is Exland's net foreign borrowing?
1
Expert's answer
2020-04-21T19:01:57-0400

a. Current Account Balance

"Current Account =(X-M)+NI+NT=700-600=100"

b. Capital Account Balance

"Capital Account=ForeignInvestments-DomesticInvestments Abroad"

"Capital Account = 350-430=-80"

c. Official Settlement Balance

"Settlement Balance= Current Account+CapitalAccount=100-80=20"

d. Net Foreign Borrowing

"Net Foreign Borrowing =Settlement Account+Holdings In ForeignCurrency=20+20=40"